Assume that a word consists of 5 lower-case letters of the alphabet. Assume that repeated letters...

  1. Assume that a word consists of 5 lower-case letters of the alphabet. Assume that repeated letters are NOT allowed to use in the word. How many words end with “t”? (4 points)
  2. Assume that a word consists of 5 lower-case letters of the alphabet. If the word needs to contain letter “o”, how many such strings are there? (4 points)
Answer #1

( a )

Number of words end with “t” of lenth 5 = 25 * 24 * 23 * 22 = 303600

( b )

Number of words contain “o” of lenth 5 = 5 * 25 * 24 * 23 * 22 = 1518000

Explanation :

(a) Ending with it MI 25 ways 23 Ways ways I way es pays en ways ho says no says I way Number of wood of length 5 ending . wi

word 25 x 24x23722 No. of S I Case 3 when letter at thind position ť present 2.5 14оу оу 24oy погу ways No. of wind - 25X24